# Fairly Accurate: Fairness-aware Multi-group Target Detection in Online Discussion

### v2

@@ 5 GAP for Multiple Targets - $GAP_{multi}$
To account for the multi-label nature of target-group detection, we propose a mathematical extension of GAP loss ($GAP_{multi}$) to handle multi-label settings. In our proposed setting, a post can target none, one, or many to all of the groups simultaneously. Our $GAP_{multi}$ loss in Eq. 7 retains the overall error (OE) term (as in Eq. 5) to preserve utility, but it augments OE with a pairwise regularization component that minimizes squared differences in cross-entropy errors between all pairs of demographic groups. This formulation avoids dependence on a global mean error (as used in Eq. 5), enabling parallelizable fairness enforcement across $\prescript{G\mkern-0.5mu}{}{C}_{2}$ group pairs. The result is a scalable and theoretically grounded loss function that directly supports fairness in complex, multi-target classification scenarios.
@@
While we could have followed the Single Objective Optimization format in Eq. 5, there are some implementation issues associated with it. Decomposing it, we can see that the Overall loss in Eq. 3 can be computed in parallel over all the groups $G$. Now one has to wait for all these parallel computations to be over, before coming to Eq. 4, the serial bottleneck component of measuring the mean error value. Only once this mean error is computed can the group-level deviations be measured. This deviation-from-mean computational approach provides a serial component acting as a bottleneck in the otherwise parallel pipeline. To bypass this serial component bottleneck and improve parallelization, we formulate our $GAP_{multi}$ in the form in Eq. 7.
$GAP_{multi}=OE+\lambda\underset{i,j\in[G],i\neq j}{\sum}\|\underbrace{CE(g=i)}_{\text{err Group i}}-\underbrace{CE(g=j)}_{\text{err Group j}}\|_{2}^{2}$ (7)
@@
The regularization term in Eq. 7 considers pairwise error differences $\|CE(g=i)-CE(g=j)\|$ for all distinct $\prescript{G\mkern-0.5mu}{}{C}_{2}$ group pairs $i,j\in[G],i\neq j$. This allows for direct comparisons between specific groups, ensuring that no group is unduly favored or marginalized relative to another. In contrast, deviation-from-mean approaches measure errors against a global aggregate (Eq. 4), which can mask disparities between individual groups. For instance, significant differences in performance between the statistical minority groups may remain undetected if their combined performance aligns with the statistical majority group’s average error. The pairwise structure enforces symmetric treatment of errors across all groups. This is critical for multi-label tasks where multiple groups can be targeted simultaneously.
$GAP_{multi}$ has $\prescript{G\mkern-0.5mu}{}{C}_{2}$ or $G(G-1)/2$ regularization terms. These pairwise error computations, despite having quadratic growth with the group cardinality $G$, are independent and can be performed in parallel. The formulation is thus computationally scalable, even for large datasets with many demographic groups. Global deviation-based approaches rely on mean computation, as serial component, therefore has bottlenecks associated with their sequential nature. In contrast, the $GAP_{multi}$ formulation allows for parallel computations of error terms over arbitrarily large $G$ without any such bottlenecks. Thus it can be integrated seamlessly into modern deep learning pipelines for efficient optimization.
Finally, the pairwise design of $GAP_{multi}$ enables the model to capture fine-grained disparities in predictive performance across specific demographic group pairs. The formulation enforces that error disparities between two targeted groups are directly addressed, enabling a more equitable treatment. This granularity is absent in deviation-from-mean approaches, which treat all groups as equally related to the mean, overlooking important group-specific dynamics of group disparities. This makes our $GAP_{multi}$ formulation well-suited for multi-label scenarios where performance disparities are not evenly distributed, providing both scalability and interpretability for fairness-sensitive applications.
5.1 Code Flow
Alg. 1 presents the logic of the overall weighted Binary Cross Entropy (WBCE) loss, where the weights $w[g]$ represent inversely scaled values of labels (0s and 1s), within each group $g$. Thus $err\_grp[g]$ equivalently maps to Balanced Accuracy (BA), an evaluation measure widely used in datasets with label imbalance. Absence of this weighting term would strictly map to standard accuracy, following the mapping from $BCE$. The overall error $err\_overall$ is then defined as summation of errors across groups. Note that we do not weight groups while adding their errors since: a) we want to treat all groups equally; and b) TensorFlow’s $bce$ function is scale independent, i.e., it produces same error value for equal ratios of mispredictions w.r.t. total samples, irrespective of sample size. For e.g., $bce$ value over 5 samples with 1 misprediction is equal to $bce$ value over 15 samples with 3 misprediction.
Algorithm 1 $overall\_loss(y\_true,y\_pred,w[g])$
@@ 5.1 Code Flow
1:Input: $y\_true,y\_pred$triangleright True and Predicted Labels
2:Input: $w[g]$triangleright Balanced weights of Group $g$
3:$y\_true\_lab=y\_true[:\,,0]$triangleright Label Info
4:$y\_true\_dem=y\_true[:\,,1]$triangleright Demographic Info
@@
6:     $pos\_grp[g]=group(y\_true\_dem==g)$triangleright Find indices
7:     $y\_true[g]=y\_true\_lab[pos\_grp[g]]$triangleright True group labels
8:     $y\_pred[g]=y\_pred[pos\_grp[g]]$triangleright Predicted group labels
9:     $err\_grp[g]=bce(y\_true[g],y\_pred[g],w[g])$triangleright wBCE
10:$err\_overall=\sum^{G}err\_grp[g]$triangleright summation of group loss
11:Output: $err\_overall$
@@
Algorithm 2 $gap\_loss(y\_true,y\_pred,w[g])$
Loss function for optimizing Group Balanced Error
0:Repeat Steps 1-10 of Alg. 1
1:for group pairs $[i,j]\in G,[i\neq j]$ do $\triangleright$ $\prescript{G\mkern-0.5mu}{}{C}_{2}$ iterations
2:     $err\_group\_pairs=\sum(err\_grp[i]-err\_grp[j])^{2}$
3:$err\_balanced=err\_overall+\lambda\cdot err\_group\_pairs$
4:Output: $err\_balanced$
@@ 6 Impossibility Results
As discussed earlier, achieving fairness in target-group detection requires accounting for both false positives (type-I errors) and false negatives (type-II errors) because both can represent harm to marginalized groups. One thus might consider Equalized Odds (EO) [21] or its surrogates [51, 7] as a viable optimizer. However, as noted in Section 2, EO is not well-suited for this task because it assumes asymmetric error costs. We now present a theoretical impossibility result demonstrating that EO and Accuracy Parity (AP) cannot generally be satisfied simultaneously unless strong assumptions (e.g., equal base rates) hold. This challenges the common misconception that EO implicitly ensures parity in group-level predictive accuracy.
The theorems are stated below, with their corresponding proofs in Appendix A. Furthermore, we highlight that in datasets with base rate disparities, enforcing EO might severely reduce the predictive accuracy for statistical minorities, at the cost of balancing error rates. For nomenclature, let $\{A\}$ be a demographic group having $P_{A}$ positive and $N_{A}$ negative samples, and let the proportion of positive to negative examples in a group be defined as $\delta_{A}^{*}=\frac{P_{A}}{N_{A}}$. For target-group detection, a positive / negative label for each group indicates that group being targeted / not targeted. Thus a false positive (or false alarm) indicates incorrectly detecting a group as being targeted, while a false negative (or miss) indicates failing to detect a group as being targeted.
@@
This theorem states that the alternative fairness measure FPNED [7] is incompatible with AP as well for most real-world datasets when base rates are unequal across groups ($\delta^{*}_{A}\neq\delta^{*}_{B}$). The constraint arises because FPNED requires equal error rates (FPR/FNR) across groups, while AP demands equal accuracy—conditions that cannot hold simultaneously under unequal base rates without trivializing the classifier (e.g., random predictions).
To understand why EO is not a desirable metric for fairness in target-group detection, consider a scenario with two marginalized demographic groups, Group-A and Group-B, that differ in how frequently they are targeted. When optimizing for EO, the model aligns true and false positive rates across both groups, ignoring the label imbalance. However, this alignment often requires redistributing errors, which disproportionately impacts the statistical minority group.
This is not ethically desirable for several reasons. First, “target group" refers to the group that a post is “directed at or about” [38], which is independent of the downstream task (e.g., whether the post is toxic). Being a frequent “target" does not inherently mean a group warrants prioritization. Second, a group that constitutes a small minority in society may be statistical minority in the data (across both toxic and non-toxic posts) simply by virtue of its size. This does not mean it should be de-prioritized.
Case Group Targeted Not Targeted TP FN FP TN Acc TPR FPR
I - Optimizing for Equalized Odds Group-A 100 100 80 20 30 70 0.75 0.80 0.30
@@
II - Optimizing for Accuracy Parity Group-A 100 100 77 23 23 77 0.77 0.77 0.23
@@
Table 2: Illustrative example of two marginalized groups with one (A) targeted more often than the other (B). Assume 200 total posts with Group-A having a higher base rate. To optimize for Equalized Odds (Case I), we set $TPR=0.80$ and $FPR=0.30$ across both groups. Alternatively, to optimize for Accuracy Parity (Case II), we set similar $Acc=0.77$ across both groups.
@@ 7 Experimental Setup
In this section, we describe the dataset used, the neural architecture, the baseline loss, and two baseline fairness losses used for comparing performance of the $GAP_{multi}$ loss. Implementation details regarding setup, data pre-processing, and related items can be found in Appendix B.
7.1 Dataset
@@
We use the Measuring Hate Speech (MHS) corpus [38] as shared in the HuggingFace library [14]. The dataset has $135,556$ posts, sourced from YouTube, Twitter and Reddit, where each post is annotated for its target demographic group(s). We select posts targeting one or more demographic-race groups, irrespective of toxicity labels (which we do not use). Annotator labels indicate whether or not each post targets one or more of seven race groups: Asian, Black, Latinx, Middle-Eastern, Native-American, Pacific-Islander and White. Figure 1(a) shows how frequently each group is targeted. Figure 1(b) shows the number of groups targeted per post [^2] For groups of size six and seven, posts target a broader range of demographic categories, often using inclusive terms like ‘non-white’, as shown in Table 1. This explains the higher frequency of such posts compared to those targeting fewer than five groups. . We take a 80%-20% stratified train-test split of the data.
(a) Number of posts targeting each group.
(b) Number of posts targeting multiple groups.
7.2 Neural Model and Baseline Measure
@@
For our neural model (Figure 2), we use DistilBERT [39] to extract feature representations for posts [^3] One can replace the DistilBERT layer with any other feature representation without altering the rest of the model.. This is followed by layers of dense neuron connections with relu activation and added biases. The dense layers are of (512, 128, 64) neurons respectively, with a dropout rate of 0.1, to prevent model overfitting. The layers end in $G$ classification nodes, each with sigmoid activation, no biases and 0.5 classification threshold. Given the multi-label binary classification setup of our problem, we use weighted Binary Cross Entropy (wBCE) as our baseline loss to optimize Overall Error (OE) for the $G$ output nodes. Our experiments freeze the weights of the DistilBERT layer, hence the only trainable parameters are the nodes in the dense layers and the final classification layer.
@@
Figure 2: Our multi target-group detection architecture. The model has shared parameters to learn both general and group-targeted language properties. The classification nodes each learn group-specific mappings.
7.3 Other Differentiable Fairness Measure
@@
CLA. We compare to CLAss-wise equal opportunity (CLA) [42], a differentiable fairness loss. CLA seeks to balance False Negative Rate (FNR) across protected groups [9, 21] by minimizing error in absolute differences between error w.r.t. a label $(BCE(y))$ vs. error w.r.t. a label given the group $(BCE(y,g))$. We replace BCE with wBCE in Eq. 8 to compare fairly under label imbalance.
$CLA=BCE+\lambda\cdot\sum_{y\in C}\sum_{g\in G}|BCE(y,g)-BCE(y)|$ (8)
@@
ADV. We also compare against a differentiable loss: ADV [50]. Since ADV only supports a 2-group setting, we present results with it in Appendix C.
7.4 Evaluation Measures Considered
@@
For model evaluation, we consider balanced accuracy, hamming loss and macro-level measures to capture a comprehensive view of model performance for various groups across losses. These measures account for group imbalances, ensuring that the statistical minority groups are fairly evaluated, while macro-level measures provide an aggregate view that treats each group equally regardless of their size.
Balanced Accuracy (BA)
Balanced Accuracy provides an assessment of accuracy while accounting for label imbalance in datasets. It computes the average accuracy of each label, thereby offering a balanced perspective to account for the unequal label distribution. By considering both the sensitivity (TPR: true positive rate) and specificity (TNR: true negative rate) of each label, BA effectively captures the model’s ability to correctly classify instances across all labels, regardless of their prevalence.
$BA=(TPR+TNR)/2.0$ (9)
Average Balanced Accuracy (Avg. BA)
When optimizing accuracy across groups, we report the macro average over the group-specific BAs as a summary statistics. The Avg. BA (macro) treats each group equally, ensuring that the classifier’s performance is evaluated in a balanced manner across all demographic groups.
$Avg.\,BA=\frac{1}{G}\sum_{g=1}^{G}BA(g)$ (10)
Hamming Loss (HL)
This is a widely employed metric for assessing the performance of multi-label classifiers. Formally, for a dataset with $N$ instances and $G$ labels, the Hamming Loss (HL) quantifies the fraction of incorrectly predicted labels across all instances in the dataset, with $hamming(y_{i}(g),\hat{y}_{i}(g))$ as an indicator function of 1 if the $g$-th label for instance $i$ is incorrectly predicted and 0 otherwise. Specifically, it measures the average fraction of labels that are misclassified in comparison to the true label set. HL and Subset Accuracy Loss are comparable under small label cases [48], hence we just report HL.
Other measures of interest
Although we are strictly optimizing for similar BA across groups, it is also important to assess the effect that this gain in fairness may have on other measures of interest, like Precision, Recall, F1. Given our multi-label setup, we report the macro variants (average over group-specific numbers) of the mentioned measures.

### v3

@@ 5 GAP for Multiple Targets - $GAP_{multi}$
Given the multi-label nature of multi-group target detection (i.e., a single post can target 0-all groups), we propose a mathematical extension of the GAP loss named $GAP_{multi}$.
@@
While we could have followed the Single Objective Optimization (SOO) format in Eq. 3, we seek to address several issues associated with it. We can see that the Overall loss in Eq. 1 can be computed in parallel over all the $G$ groups. However, Eq. 2 acts as the serial step while computing the mean error value, only after which the group-level deviations be measured. This deviation-from-mean computational approach provides a serial bottleneck in the otherwise parallel pipeline. To bypass this serial block and improve parallelization, we formulate our $GAP_{multi}$ in the form in Eq. 5.
$GAP_{multi}=OE+\lambda\underset{j,k\in G,j\neq k}{\sum}\|\underbrace{CE(g=j)}_{\text{err Group j}}-\underbrace{CE(g=k)}_{\text{err Group k}}\|_{2}^{2}$ (5)
@@
The regularization term in Eq. 5 considers pairwise error differences for all distinct $\prescript{|G|\mkern-0.5mu}{}{C}_{2}$ group pairs $j,k\in G,j\neq k$. This allows for direct comparisons between groups, ensuring that no group is unduly favored or marginalized relative to another. In contrast, deviation-from-mean approaches measure errors against a global aggregate (Eq. 2), which can mask disparities between individual pairs. For instance, significant differences in performance between the statistical minority groups may remain undetected if their combined performance aligns with the statistical majority group’s average error. The pairwise structure enforces symmetric treatment of errors across all groups. The algorithmic flow is described in details via Alg. 1 and Alg. 2 below.
Algorithm 1 overall_loss(y_true,y_pred,w[g])overall\_loss(y\_true,y\_pred,w[g])
@@
1:Input: y_true,y_predy\_true,y\_pred $\triangleright$ True and Predicted Labels
2:Input: $w[g]$ $\triangleright$ Balanced weights of Group $g$
3:y_true_lab=y_true[:,0]y\_true\_lab=y\_true[:\,,0] $\triangleright$ Label Info
4:y_true_dem=y_true[:,1]y\_true\_dem=y\_true[:\,,1] $\triangleright$ Demographic Info
@@
6:  pos_grp[g]=group(y_true_dem==g)pos\_grp[g]=group(y\_true\_dem==g) $\triangleright$ Find indices
7:  y_true[g]=y_true_lab[pos_grp[g]]y\_true[g]=y\_true\_lab[pos\_grp[g]] $\triangleright$ True group labels
8:  y_pred[g]=y_pred[pos_grp[g]]y\_pred[g]=y\_pred[pos\_grp[g]] $\triangleright$ Predicted group labels
9:  err_grp[g]=bce(y_true[g],y_pred[g],w[g])err\_grp[g]=bce(y\_true[g],y\_pred[g],w[g]) $\triangleright$ wBCE
[…]
11:Output: err_overallerr\_overall
Finally, our $GAP_{multi}$ loss retains the overall error (OE) term (as in Eq. 3) to preserve utility, but it augments OE with a pairwise regularization component enabling parallelizable fairness enforcement across $\prescript{|G|\mkern-0.5mu}{}{C}_{2}$ or |G|(|G|1)/2|G|(|G|-1)/2 group pairs. These pairwise error computations, despite quadratic growth with the group cardinality $|G|$, are independent and can be performed in parallel. The formulation is thus computationally scalable, even for large datasets with many demographic groups. Thus it can be integrated seamlessly into modern deep learning pipelines for efficient optimization.
Alg. 1 presents the logic of the overall weighted Binary Cross Entropy (WBCE) loss, where the weights $w[g]$ represent inversely scaled values of labels (0s and 1s), within each group $g$. Thus err_grp[g]err\_grp[g] equivalently maps to Balanced Accuracy (BA), an evaluation measure widely used in datasets with label imbalance. Absence of this weighting term would strictly map to standard accuracy, following the mapping from $BCE$. The overall error err_overallerr\_overall is then defined as summation of errors across groups. Note that we do not weight groups while adding their errors since: a) we want to treat all groups equally; and b) TensorFlow’s bcebce function is scale independent, i.e., it produces same error value for equal ratios of mispredictions w.r.t. total samples, irrespective of sample size. For e.g., bcebce value over 5 samples with 1 misprediction is equal to bcebce value over 15 samples with 3 misprediction.
Algorithm 2 gap_loss(y_true,y_pred,w[g])gap\_loss(y\_true,y\_pred,w[g])
Loss function for optimizing Group Accuracy Parity Error
0:Repeat Steps 1-10 of Alg. 1
1:for group pairs $j,k]\in G,[j\neq k]$ do $\triangleright$ $\prescript{|G|\mkern-0.5mu}{}{C}_{2}$ iterations
2:  $err\_group\_pairs=\sum(err\_grp[j]-err\_grp[k])^{2}$
[…]
4:Output: err_balancederr\_balanced
@@ 6 Impossibility Results
Achieving fairness in target-group detection requires accounting for both false positives (type-I errors) and false negatives (type-II errors) because both can represent harm to marginalized groups. One might consider Equalized Odds (EO) [20] or its surrogates [43, 6] as a viable optimizer. However, as noted in Section 2, EO is not well-suited for this task because it assumes asymmetric error costs. We now present a theoretical impossibility result demonstrating that EO and Accuracy Parity (AP) cannot generally be satisfied simultaneously unless strong assumptions (e.g., equal base rates) hold. This challenges the common misconception that EO implicitly ensures parity in group-level predictive accuracy.
The theorems are stated below, with their corresponding proofs in Appendix A. Furthermore, we highlight that in datasets with base rate disparities, enforcing EO might severely reduce the predictive accuracy for statistical minorities, at the cost of balancing error rates. For target-group detection, a positive / negative label for each group indicates that group being targeted / not targeted. Thus a false positive (or false alarm) indicates incorrectly detecting a group as being targeted, while a false negative (or miss) indicates failing to detect a group as being targeted.
@@
This theorem states that the alternative fairness measure FPNED [6] is incompatible with AP as well for most real-world datasets when base rates are unequal across groups ($\delta^{*}_{A}\neq\delta^{*}_{B}$). The constraint arises because FPNED requires equal error rates (FPR/FNR) across groups, while AP demands equal accuracy, conditions that cannot hold simultaneously under unequal base rates without trivializing the classifier (e.g., random predictions).
To understand why EO is not a desirable metric for fairness in target-group detection, consider a scenario with two marginalized demographic groups, Group-A and Group-B, that differ in how frequently they are targeted. When optimizing for EO, the model aligns true and false positive rates across both groups, ignoring the label imbalance. However, this alignment often requires redistributing errors, which disproportionately impacts the statistical minority group. This is not ethically desirable for several reasons. First, “target group" refers to the group that a post is “directed at or about” [35], which is independent of the downstream task (e.g., whether the post is toxic). Being a frequent “target" does not inherently mean a group warrants prioritization. Second, a group that constitutes a small minority in society may be statistical minority in the data (across both toxic and non-toxic posts) simply by virtue of its size. This does not mean it should be de-prioritized.
Table 2: Illustration of two marginalized groups with one (A) targeted more often than the other (B). Assume 200 total posts with Group-A having higher base rate. To optimize for EO (Case I), we set $TPR=0.80$ and $FPR=0.30$ across both groups. Alternatively, to optimize for AP (Case II), we set similar $Acc=0.77$ across both groups.
Group Targeted Not Targeted TP FN FP TN Acc TPR FPR
I - Optimizing for Equalized Odds
Group-A 100 100 80 20 30 70 0.75 0.80 0.30
@@
II - Optimizing for Accuracy Parity
Group-A 100 100 77 23 23 77 0.77 0.77 0.23
@@ 7 Experimental Setup
In this section, we describe the datasets used, our model architecture, the baseline loss, and two baseline fairness losses used for comparing performance of the $GAP_{multi}$ loss. Implementation details regarding setup, data pre-processing, hyperparameter tuning and related items can be found in Appendix B.
Datasets
@@
We use two datasets for evaluation: the MHS corpus [35] and HateXplain [29]. MHS is available via the HuggingFace library [13] and includes 135,556 posts sourced from YouTube, Twitter and Reddit, where each post is annotated for the demographic group(s) it targets. There are seven such groups: Asian, Black, Latinx, Middle-Eastern, Native-American, Pacific-Islander and White. We include posts targeting one or more groups, irrespective of toxicity labels (which we do not use). The Black community is the statistical majority ($\sim$22k posts), while Native American and Pacific Islander are statistical minorities ($\sim$2k posts). Details on the frequency with which different groups are targeted in the data, and the number of groups targeted by posts, are included in Appendix Figure 5. We do a random train-validation-test split of 7070-$10$-$20\%$ on the MHS corpus.
HateXplain contains 57,687 posts sourced from Twitter and Gab. There are five racial groups—African, Arab, Asian, Caucasian, and Hispanic—and the African group is the majority group ($\sim$9k posts). HateXplain comes with its own 8080-$10$-$10\%$ predefined train-validation-test split. See Appendix B for more details on the group splits across demographics and target cardinality, for both the MHS corpus and HateXplain datasets.
Incorporating data from multiple and diverse platforms is critical for validation of the different approaches compared in this work. The datasets provide multi-platform diversity, thereby allowing evaluation beyond platform-specific artifacts like slangs, structural format (e.g., lengths allowed) or distributional shifts (e.g., different communities and content having different prevalence across platforms). Both datasets contain targeted groups that are overly-represented (e.g., African//Black) and under-represented (e.g., Middle Eastern//Arab), allowing us to test the impact of sample size and class distribution on group fairness. By ensuring our proposed approach performs consistently across these varied digital signatures and distributional shifts, we demonstrate its utility for adaptable solution across content analysis tasks. Moreover, by evaluating on two different datasets, we ensure that our results do not hinge on idiosyncrasies of how any one particular dataset was sampled or annotated [16].
Model Architecture and Baseline Loss
We ran experiments using DistilBERT [36] and RoBERTa-base [27], using frozen weights to extract feature representations for posts. While RoBERTa has a richer representation than DistilBERT, in our case it did not yield any significant gain in performance, but it did add to runtime because of its larger size. Consequently, we discuss only DistilBERT results in the main body. See Appendix C for RoBERTa results.
@@
This is followed by dense neuron layers with relu activation, added biases and dropout rate of 0.1. The layers end in $|G|$ classification nodes, each with sigmoid activation, no biases and 0.5 threshold (Fig. 1). Given our multi-label binary classification setup, we use weighted Binary Cross Entropy (wBCE) as our baseline loss to optimize Overall Error (OE).
@@
Figure 1: Our multi-group target detection architecture. The model has shared parameters to learn both general and group-targeted language properties. The classification nodes each learn group-specific mappings.
Other Fairness Measures
@@
We compare to CLAss-wise equal opportunity (CLA) [38], another differentiable fairness loss that seeks to balance False Negative Rate (FNR) across groups [8, 20] by minimizing error in absolute differences between error w.r.t. a label $(BCE(y))$ vs. error w.r.t. a label given the group $(BCE(y,g))$. We replace BCE with wBCE in Eq. 6 to compare fairly under label imbalance.
$CLA=BCE+\lambda\cdot\sum_{y\in N}\sum_{g\in G}|BCE(y,g)-BCE(y)|$ (6)
@@
We also compare against another differentiable loss: ADV [42]. Since ADV only supports a 2-group setting, we present its results in Appendix D on the MHS corpus.
Evaluation Measures
@@
For model evaluation, we report Balanced Accuracy (BA) per group, Avg. BA over groups, Hamming Loss and macro-level Precision, Recall, and F1. These measures present a comprehensive view of model performance across groups that treats group equally regardless of their size. Measures are defined in Appendix C.